OpenStudy (anonymous):

Can you explain to me the steps to finding the intersection of these lines using linear combination? 2x+6y=21 -3+10y=-3

OpenStudy (anonymous):

2x+6y=21 -3x+10y=-3 ?

OpenStudy (anonymous):

I know how to do it using substitution but not combination.